Handy Hardware is a retail hardware store. Information about the stores operations follows. · November 20x1 sales amounted tThe companys balance sheet as of November 30, 20x1, is as follows: HANDY HARDWARE, INC. Balance Sheet November 30, 20x1 Asse1. Compute the budgeted cash collections for December 20x1. 2. Compute the budgeted income (loss) before income taxes for DecComplete this question by entering your answers in the tabs below. Required 1 Required 2 Required 3 Compute the budgeted incoComplete this question by entering your answers in the tabs below. Required 1 Required 2 Required 3 Compute the projected bal
Answer #1

Sales Budget December January 510000 470000 Sales Part 1 - Schedule of Expected Cash Collections December January Cash Sales27 Cash Budget 28 29 Cash Balance Beginning 30 Add Cash Collections 31 Total Cash Available 32 Less Cash Disburtments: 33 Inv

Sales Budget December 510000 January 470000 Part 1 - Schedule of Expected Cash Collections 7 December =B3+0.7 114800 =B7+B8 C27 Cash Budget 28 December 51000 =B9 =B29+B30 =B25 45900 =B33+B34+B35+B36 =B31-B37 =B40 =B40+B38 =C38 29 Cash Balance Beginni